OpenStudy (anonymous):

I am COMPLETELY lost on sequences and series stuff. How do you solve the following two: What is the sum of a 30-term arithmetic sequence where the first term is 74 and the last term is -100? What is the 22nd term of the arithmetic sequence where a1 = 8 and a9 = 56 ? Like, how do you plug those into an equation? And what equatino would you use? Just ... gah. Help x_x

OpenStudy (apoorvk):

For a sequence with 'n' terms, first term 'a' and the last term 'l', the sum 'S' is: \[S = \frac n 2 (a+l)\]

OpenStudy (apoorvk):

for the second one, form simple linear equations, to derive 'a' the first term and 'd' the common difference!

OpenStudy (anonymous):

Alright, I think I understand the first one. The second one is still a little tricky :/ What do you mean form a simple linear equation?

OpenStudy (anonymous):

wat is a1 and a9

OpenStudy (apoorvk):

For #2 a1 = 8 a9 = a1 + 8d = 56 solve these two equations to find "a1" and "d"! then, to find any nth term, use the funda- Tn = a1 + (n-1)d

OpenStudy (anonymous):

Alright! Thank you!